# Regression-Based Single-Point Zeroth-Order Optimization

### v1

@@ 4 Theoretical Convergence Analysis
Proof.
Consider any positive integer $t\geq m$. Fix any $10<c<$ 1. By our assumptions, applying Lemma 1, we find that there exists a sufficiently small absolute constant $C>0$ such that when $\eta=\frac{C}{dC_{d}L}$, we have
$\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$
for any $t\geq m$.
Next, observe that
$\ f(\bm{x}_{t+1})-f(\bm{x}_{t})$
$=$ $\ -\eta\langle\nabla f(\bm{x}_{t}),\nabla f(\bm{x}_{t})\rangle-\eta\langle\nabla f(\bm{x}_{t}),\bm{g}_{t}-\nabla f(\bm{x}_{t})\rangle+\frac{L}{2}\|\eta(\bm{g}_{t}-\nabla f(\bm{x}_{t})+\nabla f(\bm{x}_{t}))\|^{2}$
$\stackrel{{\scriptstyle\textnormal{(ii)}}}{{\mathstrut{\leq}}}$ $\ -\eta\|\nabla f(\bm{x}_{t})\|^{2}+\frac{\eta}{2}\left(\|\nabla f(\bm{x}_{t})\|^{2}+\|\xi_{t}\|^{2}\right)+\eta^{2}L\left(\|\nabla f(\bm{x}_{t})\|^{2}+\|\xi_{t}\|^{2}\right)$
$\leq$ $\ -\frac{\eta}{4}\|\nabla f(\bm{x}_{t})\|^{2}+\eta\|\xi_{t}\|^{2},$ (25)
where (4) follows by $L$-smoothness of $f$, (4) follows by application of Young’s equality as well as the notation $\xi_{t}:=\bm{g}_{t}-\nabla f(\bm{x}_{t})$, and the final inequality follows by choosing $\eta\leq\frac{1}{4L}$.
$\|\xi_{t}\|\leq\frac{1}{\sqrt{8}}\|\nabla f(\bm{x}_{t})\|,$
we find that
$\leq$ $\ -\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2},$
as desired.
Equipped with Lemma 2, we can easily show by standard arguments that the algorithm converges at a sublinear rate for smooth nonconvex functions, and at a linear rate for smooth strongly convex functions. We defer proof details to Appendix A.2.
@@
Suppose Assumptions 1 and 2 hold. Suppose $2\leq m=\mathcal{O}(d)$. Suppose we pick $\tilde{\eta}$ such that it satisfies Condition 1. Then, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, for any $T>m$, the L-RESZO algorithm achieves
$\frac{1}{T}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})|^{2}\leq\mathcal{O}\left(\frac{dC_{d}L(f(\bm{x}_{m})-f^{*}))}{T}\right).$
This convergence rate scales as $\mathcal{O}(\frac{dC_{d}L}{T})$, decaying at the rate $1/T$. When the term $C_{d}$ is on the order of $\sqrt{d}$ (as observed empirically under the adaptive choice of $\delta_{t}=\eta\|\bm{g}_{t-1}\|$ (20); see Appendix C for more details), this matches the standard convergence rate for two-point ZO of smooth nonconvex functions, up to a factor of $\sqrt{d}$.
Next, we present a linear convergence rate for strongly convex functions.
@@
Suppose Assumptions 1 and 2 hold. Suppose $2\leq m=\mathcal{O}(d)$. Suppose we pick $\tilde{\eta}$ such that it satisfies Condition 1. Then, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, for any $T>m$, the L-RESZO algorithm achieves
$f(\bm{x}_{T})-f^{*}\leq\left(1-\mathcal{O}\left(\frac{\mu}{C_{d}Ld}\right)\right)^{T-m}(f(\bm{x}_{m})-f^{*}).$
Theorem 2 shows that with the adaptive $\delta_{t}$ sequence, the L-RESZO achieves a linear convergence rate that is worse than that of two-point ZO methods for strongly convex functions only by a factor of $C_{d}$. We provide strong empirical evidence in Appendix C that $C_{d}$ scales as $\mathcal{O}(\sqrt{d})$ for both ridge and logistic regression problems. This implies that, for these problems, the theoretical linear convergence rate of L-RESZO is worse than that of two-point ZO only by a factor of $\sqrt{d}$.
@@ 5 Numerical Experiments
This section demonstrates the performance of the proposed RESZO methods in comparison with classic ZO methods through numerical experiments [^2] All numerical experiments are coded in Matlab and implemented in a computing environment with Intel(R) Core(TM) i7-1185G7 CPUs running at 3.00 GHz and with 16 GB RAM., and studies the impact of adding perturbations.
@@ 5.1 Comparison with Classic ZO Methods
$\min_{\bm{x}\in\mathbb{R}^{d}}\,f(\bm{x})=\frac{1}{2}||\bm{y}-H\bm{x}||_{2}^{2}+\frac{\lambda}{2}||\bm{x}||_{2}^{2}.$ (26)
In our experiments, each entry of the matrix $H\in\mathbb{R}^{N\times d}$ is independently drawn from the standard Gaussian distribution $\mathcal{N}(0,1)$. The column vector $\bm{y}\in\mathbb{R}^{N}$ is constructed by letting $\bm{y}=\frac{1}{2}H\bm{1}_{d}+\bm{\epsilon}$, where $\bm{1}_{d}$ denotes the $d$-dimensional all-ones vector and $\bm{\epsilon}\sim\mathcal{N}(\bm{0},0.1\bm{I})$ represents the additive Gaussian noise. The regularization parameter is fixed as $\lambda=0.1$, and the initial point is set to $\bm{x}_{0}=\bm{0}_{d}$. We set $N=1000$ and the problem dimension $d=100$.
@@
$\min_{\bm{x}\in\mathbb{R}^{d}}f(\bm{x})=\frac{1}{2}\sum_{i=1}^{N}\log\big{(}1+\exp(-y_{i}\cdot\bm{s}_{i}^{\top}\bm{x})\big{)}+\frac{\lambda}{2}||\bm{x}||_{2}^{2},$ (27)
where $\bm{s}_{i}\in\mathbb{R}^{d}$ is one of the data samples, $y_{i}\in\{-1,1\}$ is the corresponding class label, and $N=1000$ is the sample size. In our experiments, each element of a data sample $\bm{s}_{i}$ is independently drawn from the uniform distribution $\mathrm{Unif}([-1,1])$, and the corresponding label is computed as $y_{i}=\mathrm{sign}(\frac{1}{2}\bm{s}_{i}^{\top}\mathbf{1}_{d})$. The regularization parameter is fixed as $\lambda=0.1$, and we set $d=100$.
@@
which is a well-known benchmark problem for numerical optimization [25]. The global optimal solution and objective value are $\bm{x}^{*}\!=\!\bm{0}$ and $f^{*}\!=\!0$. We set $d\!=\!200$ and the initial point $\bm{x}_{0}\!=\!0.5\bm{1}_{d}$.
@@
where the decision variable is $\bm{x}\coloneq(W_{1},W_{2},W_{3},\bm{b}_{1},\bm{b}_{2},\bm{b}_{3},\bm{w}_{o})$, and $\sigma(\cdot)$ denotes the Sigmoid activation function. Let $\bm{s}_{i}\in\mathbb{R}^{n}$ and $y_{i}\in\mathbb{R}$ be the input and output of each sample $i$ with $n=6$. Thus, the problem dimension is $d=3n^{2}+4n=132$. We set the sample size to $N=500$. Each sample input $\bm{s}_{i}$ is randomly drawn from the standard Gaussian distribution $\mathcal{N}(0,1)$, and the corresponding output is computed by the neural network with ground truth $\bm{x}^{*}$, where each entry of $\bm{x}^{*}$ is randomly generated from $\mathcal{N}(0,1)$. Let the initial point be $\bm{x}_{0}\!=\!\bm{x}^{*}+\bm{\varepsilon}$, and each entry of $\bm{\varepsilon}$ is randomly drawn from the uniform distribution $\mathrm{Unif}([-1,1])$.
@@
For each ZO method, we optimize the stepsize $\eta$ and the smoothing radius $\delta$ via grid search to achieve the fastest convergence, and the selected parameters are summarized in Table 2. The same sample size $m$ is used for L-RESZO and Q-RESZO, which is set to 110, 110, 210, and 6 in Cases (a)-(d) [^3] In Cases (a)-(c), the sample size is chosen as $m=d+10$ to be slightly larger than the problem dimension $d$. However, in Case (d), a smaller sample size $m$ yields better convergence, and thus a very small sample size $m=6$ is used. It implies that, for certain problems, a small sample size might be preferable and only few recent function evaluations are needed to construct an effective gradient estimate with the proposed RESZO method., respectively. For each of the four optimization tasks described above, we run each ZO method 100 times and calculate the mean and $80\%$-confidence interval (CI) of the optimality gap $f(\bm{x})\!-\!f^{*}$. Figure 1 illustrates the convergence results using the TZO, RSZO, L-RESZO, and Q-RESZO methods, evaluated based on the number of function queries.
@@
Table 2: Selected parameters of smoothing radius $\delta$ and stepsize $\eta$ for the ZO methods.
@@
$\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$
TZO 0.002 $1.1\times 10^{-5}$ 0.01 $1.6\times 10^{-3}$ 0.01 $4.5\times 10^{-6}$ 0.01 $3.8\times 10^{-4}$
RSZO 0.2 $2.5\times 10^{-6}$ 2 $5\times 10^{-4}$ 0.5 $2\times 10^{-6}$ 0.05 $1.1\times 10^{-4}$
L-RESZO 0.002 $8\times 10^{-6}$ 0.1 $2\times 10^{-3}$ 0.02 $4.2\times 10^{-6}$ 0.001 $1.7\times 10^{-3}$
Q-RESZO 0.002 $1.6\times 10^{-5}$ 0.01 $5\times 10^{-3}$ 0.02 $1\times 10^{-5}$ 0.001 $1.7\times 10^{-3}$
As shown in Figure 1, the proposed L-RESZO and Q-RESZO methods outperform the TZO method in terms of function query complexity across all test cases and converge significantly faster than the RSZO method. Regarding iteration complexity, L-RESZO and Q-RESZO achieve a convergence rate comparable to that of TZO, which is consistent with our theoretical analysis. However, since L-RESZO and Q-RESZO require only one function query per iteration, they are approximately twice as fast as TZO with respect to function query complexity. Q-RESZO exhibits slightly faster convergence than L-RESZO, attributable to more accurate gradient estimation enabled by quadratic surrogate functions. Nevertheless, TZO ultimately attains higher convergence precision than L-RESZO and Q-RESZO. As illustrated in the next subsection, the convergence precision of L-RESZO and Q-RESZO is controlled by the smoothing radius $\delta$.
Figure 1: The convergence results of applying the TZO (4), RSZO (2), L-RESZO (Algorithm 1) and Q-RESZO (18) methods to solve ridge regression (26), logistic regression (27), Rosenbrock function minimization (28), and neural network training (29).
5.2 Impact of Random Perturbations on RESZO
Figure 2: Convergence of L-RESZO and Q-RESZO for solving the ridge regression problem (26) under different levels of perturbations and using the adaptive $\delta$ scheme (20).
To analyze the impact of perturbations added for function evaluations, we test the proposed L-RESZO and Q-RESZO methods under varying levels of the smoothing radius $\delta$ when solving the ridge regression problem (26). The experimental settings remain the same as those described in the previous subsection. Specifically, the smoothing radius $\delta$ is progressively reduced from 0.01, 0.005, 0.002, 0.001, to zero (no perturbation), and we also perform the adaptive smoothing radius scheme given by (20). The test results are illustrated in Figure 2. It is observed that a smaller smoothing radius $\delta$ leads to improved convergence precision. That is because the RESZO method essentially estimates the gradient at the perturbed point, i.e., $\nabla f(\hat{\bm{x}})=\nabla f(\bm{x}+\delta\bm{u}_{t})$, and thus a smaller $\delta$ results in lower estimation errors. However, the introduction of random perturbations is necessary for the stability of RESZO. Figure 2 shows that, when no perturbation is added ($\delta=0$), the Q-RESZO method exhibits large oscillations and poor convergence precision, while the L-RESZO method even diverges. Moreover, corroborating our theoretical results, our adaptive $\delta$ scheme given by (20) can effectively enhance convergence precision by adjusting $\delta$ based on the updated gradient estimate.
6 Conclusion
In this paper, we propose a novel yet simple Single-point Zeroth-order Optimization (SZO) framework, termed Regression-Based SZO (RESZO), which significantly improves gradient estimation by leveraging historical function evaluations from previous iterations. We theoretically establish that the RESZO method achieves convergence rates comparable to those of the two-point ZO method for both smooth nonconvex and strongly convex functions. Moreover, extensive numerical experiments demonstrate that RESZO not only matches but achieves approximately twice as fast convergence as the two-point ZO method in terms of function query complexity in all test cases. To the best of our knowledge, this is the first work to present single-point ZO that achieves a comparable convergence rate to two-point ZO and empirically outperforms it.
Limitations. Despite showing comparable convergence rates to two-point ZO, the theoretical analysis in this paper can be further enhanced by (i) rigorously bounding the term $C_{d}$ in Assumption 2 and (ii) improving the convergence rate bounds to exactly match those of two-point ZO, as observed in practice. Additionally, this work focuses exclusively on the noiseless setting; extending the analysis and algorithms to handle noisy function evaluations would better reflect practical scenarios. Future studies can be conducted to explore various strategies to further improve the RESZO framework, such as fitting more advanced surrogate functions, incorporating weighting schemes for historical function evaluations, adopting adaptive stepsize rules, and integrating trust region methods, etc.
Appendix A Proofs for Theoretical Convergence Analysis
A.1 More on Assumption 2

### v2

@@ 4 Theoretical Convergence Analysis
With Lemma 2 in place, we can show by standard arguments that the L-ReSZO algorithm converges at a sublinear rate for smooth nonconvex functions and at a linear rate for smooth strongly convex functions, as stated in Theorems 1 and 2. We defer the detailed proof to Appendix B.
@@
Suppose that Assumptions 1 and 2 hold, and that $d<m=\mathcal{O}(d)$. Suppose that $\tilde{\eta}$ is picked such that Condition 1 is satisfied. Then, by choosing $\eta=C\frac{1}{mC_{d}L}$ for some sufficiently small absolute constant $C>0$, for any $T>m$, the L-ReSZO algorithm achieves
$\frac{1}{T}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})||^{2}\leq\mathcal{O}\left(\frac{dC_{d}L(f(\bm{x}_{m})-f^{*})}{T}\right).$ (25)
This convergence rate scales as $\mathcal{O}(\frac{dC_{d}L}{T})$ when we pick $m=\Theta(d)$, decaying at the rate $1/T$. We briefly comment on picking $m$ here. Our theory suggests that $m$ should be larger than $d$ to ensure that the matrix $S_{t}$ is full-rank, a requirement for our analysis of Assumption 2); while it should not exceed $\Theta(d)$, since $\eta$ scales as $\mathcal{O}(1/m)$ and increasing $m$ slows down the convergence rate.
Next, we present the linear convergence result for strongly convex functions.
@@
Suppose that Assumptions 1 and 2 hold, and that $d<m=\mathcal{O}(d)$. Suppose that $\tilde{\eta}$ is picked such that Condition 1 is satisfied. Then, by choosing $\eta=C\frac{1}{mC_{d}L}$ for some sufficiently small absolute constant $C>0$, for any $T>m$, the L-ReSZO algorithm achieves
$f(\bm{x}_{T})-f^{*}\leq\left(1-\mathcal{O}\big(\frac{\mu}{C_{d}Ld}\big)\right)^{T-m}(f(\bm{x}_{m})-f^{*}).$ (26)
Theorem 2 shows that the L-ReSZO algorithm achieves a linear convergence rate for strongly convex functions, which is slower than that of two-point ZO methods only by a factor of $C_{d}$.
@@ 5 Numerical Experiments
This section demonstrates the performance of the proposed ReSZO methods in comparison with classic ZO methods through numerical experiments across various optimization problems, and studies the impact of adding perturbations.
@@ 5.1 Comparison with Classic ZO Methods
[…]
In our experiments, each entry of the matrix $H\in\mathbb{R}^{N\times d}$ is independently drawn from the standard Gaussian distribution $\mathcal{N}(0,1)$. The column vector $\bm{y}\in\mathbb{R}^{N}$ is constructed by letting $\bm{y}=\frac{1}{2}H\bm{1}_{d}+\bm{\epsilon}$, where $\bm{1}_{d}$ denotes the $d$-dimensional all-ones vector and $\bm{\epsilon}\sim\mathcal{N}(\bm{0},0.1\bm{I})$ represents the additive Gaussian noise. The regularization parameter is fixed as $\lambda=0.1$, and the initial point is set to $\bm{x}_{0}=\bm{0}_{d}$. We set $N=1500$ and the problem dimension $d=500$.
@@
[…]
where $\bm{s}_{i}\in\mathbb{R}^{d}$ is one of the data samples, and $y_{i}\in\{-1,1\}$ is the corresponding class label. Each element of a data sample $\bm{s}_{i}$ is independently drawn from the uniform distribution $\mathrm{Unif}([-1,1])$, and the corresponding label is computed as $y_{i}=\mathrm{sign}(\frac{1}{2}\bm{s}_{i}^{\top}\mathbf{1}_{d})$. The regularization parameter is fixed as $\lambda=0.1$. We set $N=1000$ and $d=100$.
@@
which is a well-known benchmark problem for numerical optimization [23]. The global optimal solution and objective value are $\bm{x}^{*}\!=\!\bm{0}$ and $f^{*}\!=\!0$. We set d=200d\!=\!200 and the initial point $\bm{x}_{0}\!=\!0.5\bm{1}_{d}$.
@@
where $\bm{x}\coloneq(W_{1},W_{2},W_{3},\bm{b}_{1},\bm{b}_{2},\bm{b}_{3},\bm{w}_{o})$, and $\sigma(\cdot)$ denotes the Sigmoid activation function. Let $\bm{s}_{i}\in\mathbb{R}^{n}$ and $y_{i}\in\mathbb{R}$ be the input and output of each sample $i$ with $n=6$. Thus, the problem dimension is $d=3n^{2}+4n=132$. We set the sample size to $N=500$. Each sample input $\bm{s}_{i}$ is randomly drawn from the standard Gaussian distribution $\mathcal{N}(0,1)$, and the corresponding output is computed by the neural network with ground truth $\bm{x}^{*}$, where each entry of $\bm{x}^{*}$ is randomly generated from $\mathcal{N}(0,1)$. Let the initial point be $\bm{x}_{0}\!=\!\bm{x}^{*}+\bm{\varepsilon}$, and each entry of $\bm{\varepsilon}$ is randomly drawn from the uniform distribution $\mathrm{Unif}([-1,1])$.
@@
We optimize the stepsize $\eta$ for each ZO method and the smoothing radius $\delta$ for RSZO and TZO via grid search to achieve the fastest convergence. The adaptive smoothing radius scheme (10) is employed in L-ReSZO and Q-ReSZO. The selected parameters of $\eta$ and $\delta$, as well as the sample size $m$ for regression, are presented in Table 2 in Appendix A. For each of the four optimization tasks described above, we run each ZO method 50 times and calculate the mean and $80\%$-confidence interval (CI) of the optimality gap $f(\bm{x})\!-\!f^{*}$. Figure 1 illustrates the convergence results using the TZO, RSZO, L-ReSZO, and Q-ReSZO methods, evaluated based on the number of function queries.
Figure 1: The convergence and 80% confidence intervals (CI) of TZO (4), RSZO (2), L-ReSZO (Algorithm 1) and Q-ReSZO (19) for solving ridge regression (27), logistic regression (28), Rosenbrock function minimization (29), and neural network training (30).
As shown in Figure 1, the proposed L-ReSZO and Q-ReSZO methods converge two to three times faster than TZO in terms of function query complexity and significantly outperform RSZO across all test cases. With respect to iteration complexity, L-ReSZO and Q-ReSZO achieve convergence rates comparable to, or slightly faster than, TZO. Since L-ReSZO and Q-ReSZO require only one function query per iteration, they are approximately two to three times more efficient than TZO in terms of function query complexity. In addition, Q-ReSZO exhibits slightly faster convergence than L-ReSZO, which can be attributed to its more accurate gradient estimation enabled by quadratic surrogate functions.
5.2 Effect of Perturbations on ReSZO
To analyze the effect of perturbations used in function evaluations, we evaluate the proposed L-ReSZO and Q-ReSZO methods under varying levels of the smoothing radius $\delta$ when solving the ridge regression problem (27). The experimental setup remains the same as that described in the previous subsection. Specifically, $\delta$ is progressively reduced from 1, 0.1, 0.01, 0.001, to zero (no perturbation), and we also perform the adaptive smoothing radius scheme (10).
The test results are illustrated in Figure 2. It is observed that a smaller smoothing radius $\delta$ generally leads to improved convergence precision. This is because ZO methods estimate the gradient at a perturbed point, i.e., $\nabla f(\hat{\bm{x}})=\nabla f(\bm{x}+\delta\bm{u}_{t})$, and thus a smaller $\delta$ results in lower gradient estimation errors. However, the introduction of random perturbations is necessary for the stability of ReSZO. As shown in Figure 2, when no perturbation is applied ($\delta=0$) or when the perturbation is excessively small ($\delta=0.001$), the L-ReSZO method diverges, and the Q-ReSZO method exhibits large oscillations in the former case. The proposed adaptive smoothing radius scheme effectively addresses these issues by automatically adjusting $\delta$, achieving both algorithmic stability and high convergence precision.
Figure 2: Convergence of L-ReSZO and Q-ReSZO for solving the ridge regression problem (27) under different levels of perturbations and using the adaptive smoothing radius scheme (10).
6 Conclusion
This paper proposes a novel yet simple single-point zeroth-order optimization (SZO) framework termed Regression-based SZO (ReSZO), which significantly improves gradient estimation by leveraging historical function evaluations from previous iterations. We theoretically establish that the ReSZO method achieves convergence rates comparable to those of the two-point ZO method for both smooth nonconvex and strongly convex functions. Furthermore, extensive numerical experiments demonstrate that ReSZO not only matches but empirically converges approximately two to three times faster than two-point ZO methods in terms of function query complexity across all test cases. To our knowledge, this is the first single-point zeroth-order method that attains convergence rates comparable to two-point methods while achieving superior empirical performance. Future work will explore strategies to further enhance the ReSZO framework, including more expressive surrogate models, weighted regression schemes, and extensions to noisy function evaluations that better reflect practical settings.
Appendix A Numerical Experiment Settings
For the numerical experiments in Section 5.1, we optimize the stepsize $\eta$ for each ZO method and the smoothing radius $\delta$ for RSZO and TZO via grid search to achieve the fastest convergence. The adaptive smoothing radius scheme (10) is employed in L-ReSZO and Q-ReSZO. The selected parameters of $\eta$ and $\delta$, as well as the sample size $m$ for regression, are presented in Table 2.
The same sample size $m$ is used for both L-ReSZO and Q-ReSZO, with $m$ set to 510, 110, 210, and 6 for Cases (a)–(d), respectively. In Cases (a)–(c), the sample size is chosen as $m=d+10$, slightly exceeding the problem dimension $d$. In Case (d), a smaller sample size yields better convergence, and thus $m=6$ is adopted. This observation suggests that, for certain problems, only a small number of recent function evaluations is sufficient to construct an accurate gradient estimate using the proposed ReSZO method.
@@ Appendix A Numerical Experiment Settings
Table 2: The selected smoothing radius $\delta$, stepsize $\eta$, and sample size $m$ for the ZO methods in Cases (a)-(d) in Section 5.1.
@@
$\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{m}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{m}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{m}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{m}$
TZO 0.01 $1.2\times 10^{-6}$ 0.01 $1.6\times 10^{-3}$ 0.01 $4.5\times 10^{-6}$ 0.01 $3.8\times 10^{-4}$
RSZO 0.2 $3\times 10^{-7}$ 2 $5\times 10^{-4}$ 0.5 $2\times 10^{-6}$ 0.05 $1.1\times 10^{-4}$
L-ReSZO adaptive $1.5\times 10^{-6}$ 510 adaptive $8\times 10^{-3} 11$0 adaptive $4.2\times 10^{-6}$ 210 adaptive $1.7\times 10^{-3}$ 6
Q-ReSZO adaptive $2\times 10^{-6}$ 510 adaptive $5\times 10^{-3}$ 110 adaptive $1\times 10^{-5}$ 210 adaptive $1.8\times 10^{-3}$ 6
Appendix B Proofs for Theoretical Convergence Analysis
B.1 Theoretical justification for Assumption 2